A metallic sphere floats in an immiscible mixture of water (ρw = 103 kg/m3) & a liquid (ρl = 13.5 × 103 kg/m3 ) such that its 4/5 th volume is in water & its 1/5 th volume in the liquid. Find the density of metal

Grade:Upto college level

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Mainak


let the density of metal is ρ  ,and volume of the metal is V

ρVg = ρw(4V/5)g  + ρl(V/5)g

 ρ = ρw(4/5)  + ρl(1/5)

   =3.5 × 103 kg/m3

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin